# Aspects of the commuting graph [^1] To the memory of Richard Parker

### v3

@@ Abstract
The commuting graph of a group $G$ is the graph whose vertices are the elements of $G$, two distinct vertices joined if they commute.
@@ 1 Introduction
Graphs associated with some algebraic structures have been actively investigated, such as, the power graph of a group (cf. [1, 20, 22, 24]), and of course the famous Cayley graphs which have a long history. Moreover, graphs from algebraic structures have valuable applications (cf. [45]): for example, Cayley graphs as classifiers for data mining [46].
Let $G$ be a group and $S$ be a subset of $G$. The commuting graph of $G$ on $S$ is a simple graph with vertex set $S$, and two distinct elements are adjacent in this graph if they commute in $G$. In particular, if $S=G$, then we denote this commuting graph on $G$ by $\Gamma(G)$. In 1955, Brauer and Fowler [12] first introduced the commuting graph of a group and showed that there are only finitely many simple groups with even order such that they have a prescribed centralizer. In fact, in the seminal paper [12], they considered the group commuting graph on the set of all non-trivial elements. Since then, the commuting graphs have received considerable attention and were well studied in the literature. For example, Segev et al. [64, 65, 66] established some longstanding conjectures on division algebras by using some combinatorial parameters of commuting graphs. Britnell and Gill [15] classified all finite quasi-simple groups whose commuting graph is perfect, where the vertex set of the commuting graph is the set consisting of all non-central elements. Giudici and Parker [38] showed that for every positive integer $d$, there is a finite $2$-group $G$ such that the commuting graph of $G$ on the set of all non-central elements has diameter greater than $d$; further results are found in [30]. Morgan and Parker [58] proved that, for a finite group $G$ with trivial center, every connected component of the commuting graph of $G$ on the set of all non-trivial elements has diameter at most $10$. Recently, Lewis et al. [9] extended some results by Morgan and Parker [58] about commuting graphs. Moreover, some further applications of commuting graphs in group theory can be found in [21].
The first part of this paper is devoted to the algorithmic question: Given a graph $\Gamma$, is it isomorphic to the commuting graph of a group? We give a quasi-polynomial algorithm in general, and a polynomial algorithm in the case of extraspecial $p$-groups with $p$ odd.
A number of important graph classes, including perfect graphs, cographs, chordal graphs, split graphs, and threshold graphs, can be defined either structurally or in terms of forbidden induced subgraphs. Forbidden subgraphs of power graphs were studied in [54]. In [21], Cameron described various graphs whose vertex set is the universe of a group and whose edges reflect group structure in some way, and proposed the following question.
@@
Britnell and Gill [15] studied the finite groups whose commuting graph is perfect. Ma and Cameron [51] completely classified the finite groups whose commuting graph is a split graph or a threshold graph.
In the second part of the paper, we address the question of when the commuting graph is a cograph or a chordal graph. We prove that in some sense, the class of such groups generalizes the class of finite solvable groups (see Theorem 11.7 in Subsection 11.3). Also we determine all finite simple groups whose commuting graph is a cograph (see Problem 1 in Subsection 11.3 and Theorem 12.15 at the end of Section 12). The results partly answer Question 1.
These issues are relevant to the recognition question. There are extra hypotheses which could make the problem of recognition of commuting graphs easier. One is to restrict the groups to a known class, such as nilpotent groups. Another is to restrict the graphs to a class in which computation is easier. For example, the result of the second and third authors describes precisely the split or threshold graphs which are commuting graphs; these are the complete graphs, together with the graphs consisting of an $n$-vertex complete graph (where $n$ is odd), with $n$ pendant edges added at one of its vertices. Clearly these are easily recognised. We have not pursued this issue further.
We note that some works, especially on connectedness and diameter, define the commuting graph to have vertex set $G\setminus Z(G)$ (that is, omitting dominating vertices). This has little effect on our results. For the recognition problem, if we are given the commuting graph on $G\setminus Z(G)$, then the commuting graph on $G$ is found by adding $|Z(G)|$ dominating vertices to the graph, and this number is a divisor of $|G\setminus Z(G)|$ so there are not many possibilities. For the second part of the paper, see Proposition 10.2.
We conclude this section with a suggestion for further research, which would link the two parts of the paper together more closely.
Question 2
@@ 2 Preliminaries
Every group considered is finite. Throughout, $G$ always denotes a finite group and $e_{G}$ or $1_{G}$ denotes the identity element of $G$, and we denote $e_{G}$ simply by $e$ or $1$ if the considered situation is unambiguous. Let $g\in G$. The order $o(g)$ of $g$ in $G$ is the size of the cyclic subgroup $\langle g\rangle$ generated by $g$. The element $g$ is called an involution if $o(g)=2$. The center of $G$, denoted by $Z(G)$, is defined as
@@
Also, the centralizer $C_{G}(g)$ of $g$ in $G$ is the set of all elements that commute with $g$, that is,
@@
Note that $C_{G}(g)$ is a subgroup of $G$. The commutator of two elements $x,y$ is denoted as $[x,y]$, that is, $[x,y]=xyx^{-1}y^{-1}$. Note that $x$ and $y$ commute if and only if $[x,y]=e$. Denote by $\mathbb{Z}_{n}$ the cyclic group with order $n$. For a prime $p$, an elementary abelian $p$-group is a direct product of cyclic groups of order $p$. A Sylow $p$-subgroup of $G$ is a $p$-subgroup whose index is coprime to $p$.
If $G$ is a finite group, the following conditions are equivalent:
@@
    $G$ is nilpotent;
@@
    every Sylow subgroup of $G$ is normal in $G$;
@@
Given a positive integer $n$, we use $D_{2n}$ to denote the dihedral group with order $2n$, a presentation of $D_{2n}$ is given by
$D_{2n}=\langle a,b:a^{n}=b^{2}=e,bab=a^{-1}\rangle.$ (1)
@@
Suppose that $A$ is an abelian group. The generalized dihedral group corresponding to $A$, denoted by $D(A)$, is the external semidirect product of $A$ with the cyclic group of order two, with the non-identity element acting as the inverse map on $A$. Viewing this external semidirect product as an internal semidirect product, $A$ is an abelian normal subgroup of index two. A presentation for $D(A)$ is as follows:
$D(A)=\langle A,x:o(x)=2,xgx=g^{-1}\hbox{ for each $g\in A$}\rangle.$ (3)
@@
Given a positive integer $m$ at least $2$, Johnson [44, pp. 44–45] introduced the generalized quaternion group $Q_{4m}$ of order $4m$, that is,
$Q_{4m}=\langle x,y:x^{m}=y^{2},x^{2m}=e,y^{-1}xy=x^{-1}\rangle.$ (4)
Note that for any $1\leq i\leq 2m$, $o(x^{i}y)=4$ and $Z(Q_{4m})=\{e,x^{m}\}$.
@@
be the Frobenius kernel of $G$. It is well-known (see, for example, [3, 35.24 and 35.25]) that $K\trianglelefteq G$, $G=K\rtimes H$, $C_{G}(h)\leq H$ for each $h\in H$, and $C_{G}(k)\leq K$ for each $k\in K$. Moreover, by the Thompson theorem on finite groups with fixed-point-free automorphisms of prime order [74, Theorem 1], $K$ is nilpotent. Frobenius groups are also characterised as transitive non-regular permutation groups in which non-identity elements fix at most one point.
Every graph considered is simple: that is, it is undirected, with no loops and no multiple edges. Given a graph $\Gamma$, the vertex set and edge set of $\Gamma$ are denoted by $V(\Gamma)$ and $E(\Gamma)$, respectively. Let $\mathcal{F}$ be a collection of finite graphs. If no induced subgraph of $\Gamma$ is isomorphic to a graph in $\mathcal{F}$, then $\Gamma$ is called $\mathcal{F}$-free; and if =$\mathcal{F}=\{\Delta\}$, we say that $\Gamma$ is $\Delta$-free. Sometimes we say instead that $\Gamma$ forbids $\mathcal{F}$ or $\Delta$. Denote by $P_{n}$ and $C_{n}$ the path of length $n-1$ and the cycle of length $n$, respectively. For distinct $x,y\in V(\Gamma)$, if $x$ is adjacent to $y$ in $\Gamma$, then we use $x\sim_{\Gamma}y$ to denote this edge $\{x,y\}$, shortly by $x\sim y$. In particular, denote by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}$
the induced path of length $n-1$, which is a subgraph of $\Gamma$ induced by the set $\{x_{1},x_{2},\ldots,x_{n}\}$, and is isomorphic to $P_{n}$ with $\{x_{i},x_{i+1}\}\in E(\Gamma)$ for all $11\leq i\leq n-$ 1. Similarly, in $\Gamma$, the induced cycle of length $n$ by $\{x_{1},x_{2},\ldots,x_{n}\}$ is an induced subgraph, and is denoted by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}\sim x_{1},$
where $\{x_{1},x_{n}\}\in E(\Gamma)$ and $\{x_{i},x_{i+1}\}\in E(\Gamma)$ for all $11\leq i\leq n-$ 1.

### v4

@@ Abstract
The commuting graph of a group $G$ is the simple graph whose vertices are the elements of $G$, two distinct vertices joined if they commute.
@@ 1 Introduction
Graphs associated with some algebraic structures, such as the power graph of a group or semigroup, have been actively investigated (cf. [1, 20, 22, 24]) Of course the famous Cayley graphs have a long history. Moreover, graphs from algebraic structures have valuable applications (cf. [47]): for example, Cayley graphs as classifiers for data mining [48].
Let $G$ be a group and $S$ be a subset of $G$. The commuting graph of $G$ on $S$ is a simple graph with vertex set $S$, and two distinct elements are adjacent in this graph if they commute in $G$. In particular, if $S=G$, then we denote this commuting graph on $G$ by $\Gamma(G)$. In 1955, Brauer and Fowler [12] first introduced the commuting graph of a group and showed that there are only finitely many simple groups with even order such that they have a prescribed centralizer. In fact, in the seminal paper [12], they considered the induced subgraph of the commuting graph on the set of all non-trivial elements of the group. Since then, the commuting graphs have received considerable attention and have been well studied in the literature. For example, Segev et al. [66, 67, 68] established some longstanding conjectures on division algebras by using some combinatorial parameters of commuting graphs. Britnell and Gill [15] classified all finite quasi-simple groups whose commuting graph is perfect, where the vertex set of the commuting graph is the set consisting of all non-central elements. Giudici and Parker [39] showed that for every positive integer $d$, there is a finite $2$-group $G$ such that the commuting graph of $G$ on the set of all non-central elements has diameter greater than $d$; further results are found in [30]. Morgan and Parker [60] proved that, for a finite group $G$ with trivial centre, every connected component of the commuting graph of $G$ on the set of all non-trivial elements has diameter at most $10$. Recently, Lewis et al. [9] extended some results by Morgan and Parker [60] about commuting graphs. Moreover, some further applications of commuting graphs in group theory can be found in [21].
The first part of this paper is devoted to the following algorithmic question.
@@
Given a graph $\Gamma$, is it isomorphic to the commuting graph of a group?
We give a quasi-polynomial algorithm in general, and a polynomial-time algorithm in the case of extraspecial $p$-groups with $p$ odd.
A number of important graph classes, including perfect graphs, cographs, chordal graphs, split graphs, and threshold graphs, can be defined either structurally or in terms of forbidden induced subgraphs. Forbidden subgraphs of power graphs were studied in [56]. In [21], the second author described various graphs that can be associated with a given group $G$. The vertex set of such a graph is the set of all elements of $G$, and its edge set reflects the group structure of $G$ in some way. He proposed the following question.
Question 2
@@
Britnell and Gill [15] studied the finite groups whose commuting graph is perfect. The second and third authors [53] completely classified the finite groups whose commuting graph is a split graph or a threshold graph.
In Part II of the paper (Sections 10, 11, 12, and 13) , we address the question of when the commuting graph is a cograph or a chordal graph. We prove that in some sense, the class of such groups generalizes the class of finite solvable groups (see Theorem 11.7 in Subsection 11.3). Also we determine all finite simple groups whose commuting graph is a cograph (see Problem 1 in Subsection 11.3 and Theorem 12.15 at the end of Section 12). The results partly answer Question 2.
These issues are relevant to the recognition question (Question 1). There are extra hypotheses which could make the problem of recognition of commuting graphs easier and yield faster algorithms. One is to restrict the groups to a known class, such as nilpotent groups. Another is to restrict the graphs to a class in which computation is easier. For example, the result of the second and third authors [53] describes precisely the split or threshold graphs which are commuting graphs; these are the complete graphs, together with the graphs consisting of an $n$-vertex complete graph (where $n$ is odd), with $n$ pendant edges added at one of its vertices. Clearly these are easily recognised. We have not pursued this issue further.
We note that some works, especially on connectedness and diameter, define the commuting graph to have vertex set $G\setminus Z(G)$ (that is, omitting dominating vertices). This has little effect on our results. For the recognition problem, if we are given the commuting graph on $G\setminus Z(G)$, then the commuting graph on $G$ is found by adding $|Z(G)|$ dominating vertices to the graph, and this number is a divisor of $|G\setminus Z(G)|$ so there are not many possibilities. In the second part of the paper too, as mentioned before Proposition 10.2, we can define the commuting graph without including the centre $Z(G)$ and the results essentially remain the same.
We conclude the introduction with a suggestion for further research, which would link the two parts of the paper together more closely.
Question 3
@@ 2 Preliminaries
Every group considered is finite. Throughout, $G$ always denotes a finite group and $1_{G}$ denotes the identity element of $G$, and we denote $1_{G}$ simply by $1$ if the considered group is unambiguous. The order of the group $G$ is its cardinality, denoted $|G|$. For an element $g\in G$, the order $o(g)$ of $g$ in $G$ is the size of the cyclic subgroup $\langle g\rangle$ generated by $g$. The element $g$ is called an involution if $o(g)=2$. The centre of $G$, denoted by $Z(G)$, is defined as the subgroup
@@
Also, the centralizer $C_{G}(g)$ of $g$ in $G$ is the subgroup of all elements that commute with $g$, that is,
@@
The commutator of two elements $x,y$ is denoted as $[x,y]$, that is, $[x,y]=x^{-1}y^{-1}xy$. Note that $x$ and $y$ commute if and only if $[x,y]=1$. The derived group, or commutator subgroup, $G^{\prime}$ of $G$ is the subgroup generated by all commutators. The derived series is the series of subgroups $G^{(n)}$ of $G$ defined by $G^{0}=G$ and $G^{(n+1)}=(G^{(n)})^{\prime}$. The group $G$ is solvable of $G^{(n)}=1$ for some $n$.
Denote by $\mathbb{Z}_{n}$ the cyclic group with order $n$. For a prime $p$, a $p$-group is a group whose order is a power of $p$. An elementary abelian $p$-group is a $p$-group that is a direct product of cyclic groups of order $p$. A Sylow $p$-subgroup of $G$ is a $p$-subgroup whose index in $G$ is coprime to $p$.
An automorphism of $G$ is a permutation $\alpha$ of $G$ preserving the group operation: $(gh)^{\alpha}=g^{\alpha}h^{\alpha}$. An inner automorphism is a map $\phi(g):x\mapsto g^{-1}xg$. It is readily checked that the automorphisms of $G$ form a group $\mathop{\mathrm{Aut}}(G)$, while the inner automorphisms form a subgroup $\mathop{\mathrm{Inn}}(G)$ of $\mathop{\mathrm{Aut}}(G)$.
A subgroup $N$ of $G$ is a normal subgroup of $G$, denoted $N\lhd G$, if $gN=Ng$ for all $g\in G$. The quotient $G/N$ is the group whose elements are the cosets of $N$ in $G$ with the (well-defined) group operation $(Ng_{1})(Ng_{2})=N(g_{1}g_{2})$. For example,
@@
    the centre $Z(G)$ is a normal subgroup of $G$, and $G/Z(G)$ is isomorphic to $\mathop{\mathrm{Inn}}(G)$;
@@
    the inner automorphism group $\mathop{\mathrm{Inn}}(G)$ is a normal subgroup of the automorphism group $\mathop{\mathrm{Aut}}(G)$; the quotient is the outer automorphism group $\mathop{\mathrm{Out}}(G)$ of $G$;
@@
    the derived group $G^{\prime}$ is a normal subgroup of $G$; it is the smallest normal subgroup $N$ for which $G/N$ is abelian.
A subgroup of $G$ is characteristic if it is mapped to itself by all automorphisms of $G$. Since a subgroup is normal if it is mapped to itself by all inner automorphisms, we see that each characteristic subgroup is normal.
A group $G$ is simple if its only normal subgroups are $G$ and $\{1\}$. In this paper, we will use the Classification of the Finite Simple Groups (CFSG), which will be explained when it is used in Section 8.1. Here we note one consequence of CFSG, the proof of the Schreier conjecture:
Theorem 2.1
If $G$ is a non-abelian finite simple group, then $\mathop{\mathrm{Out}}(G)$ is solvable.
Each minimal (non-trivial) normal subgroup of a finite group is a direct product of isomorphic simple groups [32, Theorem 4.3A(iii)]. The socle of $G$, denoted $\mathop{\mathrm{Soc}}(G)$, is the product of the minimal normal subgroups of $G$; it is a direct product of simple groups (not necessarily isomorphic). If the socle is simple, we say that $G$ is almost simple.
A normal series for a group $G$ is a collection of normal subgroups $N_{i}\lhd G,0\leq i\leq r$ such that
$\{1\}=N_{0}\subset N_{1}\subset N_{2}\subset\cdots\subset N_{r}=G.$
This series is a central series for $G$ if $N_{i}/N_{i-1}\subseteq Z(G/N_{i-1})$ for $1\leq i\leq r$. A group $G$ is nilpotent if it has a central series. If $G$ is nilpotent, and we build the upper central series by taking $N_{i}/N_{i-1}=Z(G/N_{i-1})$, the number of steps in the series is the nilpotence class of $G$.
Theorem 2.2
[45, Theorem 1.26] If $G$ is a finite group, the following conditions are equivalent:
  1. (a)
    $G$ is nilpotent.
  2. (b)
    Every Sylow subgroup of $G$ is normal in $G$.
  3. (c)
@@
A component of $G$ is a subgroup $H$ of $G$ which is quasisimple (that is, $H=H^{\prime}$ and $H/Z(H)$ is simple) and subnormal in $G$ (that is, a term in some normal series of $G$). Let $\mathop{\mathrm{Comp}}(G)$ be the set of components of $G$. Then the subgroup $E(G)$ is the subgroup generated by $\mathop{\mathrm{Comp}}(G)$, and the generalized Fitting subgroup $F^{*}(G)$ of $G$ is the central product
$F^{*}(G)=F(G)\circ E(G),$ (1)
see [3, 31.12].
Given a positive integer $n$, we use $D_{2n}$ to denote the dihedral group with order $2n$. A presentation of $D_{2n}$ is given by
$D_{2n}=\langle a,b:a^{n}=b^{2}=1,bab=a^{-1}\rangle.$ (2)
@@
Suppose that $A$ is an abelian group. The generalized dihedral group corresponding to $A$, denoted by $D(A)$, is the external semidirect product $A\rtimes T$ of $A$ with the cyclic group $T=\langle t\rangle$ of order two, with the non-identity element $t$ acting as the inverse map on $A$, that is,
$x^{t}=x^{-1}\textrm{ for all }x\in A.$
Viewing this as an internal semidirect product, $A$ is an abelian normal subgroup of $D(A)$ of index two. A presentation for $D(A)$ is as follows:
$D(A)=\langle A,t:o(t)=2,tgt=g^{-1}\hbox{ for each $g\in A$}\rangle.$ (4)
@@
Proof 1
As $A$ is a normal subgroup of $D(A)$ of index two, every element $g\in D(A)\setminus A$ is in the coset $At$ and is of the form $g=at$. Hence, $g^{2}=atat=aa^{t}=aa^{-1}=1$ which means $g$ is an involution. Furthermore, $D(A)$ is commutative if and only if for all $a\in A$, $at=ta$. That is, $tat=a^{-1}=a$ which implies $a^{2}=1$ for all $a\in A$, implying that $A$ is an elementary abelian $2$-group which, in turn, implies that $D(A)$ is an elementary abelian $2$-group. The last part follows directly from the definition of dihedral groups. $\qed$
@@
Given a positive integer $m$ at least $2$, Johnson [46, pp. 44–45] introduced the generalized quaternion group $Q_{4m}$ of order $4m$, that is,
$Q_{4m}=\langle x,y:x^{m}=y^{2},x^{2m}=1,y^{-1}xy=x^{-1}\rangle.$ (5)
Observation 2
For any $1\leq i\leq 2m$, $o(x^{i}y)=4$ and $Z(Q_{4m})=\{1,x^{m}\}$.
Proof 2
By definition, $x^{2m}=y^{4}=1$, which implies that $o(y)=4$, as there are no other relations. Furthermore, as $xy=yx^{-1}$ we can write $x^{i}y=yx^{-i}$ for any positive integer $i$. It follows that $x^{i}yx^{i}y=yx^{-i}x^{i}y=y^{2}$ for all $i$. Hence, $(x^{i}y)^{4}=y^{4}=1$ for all $i$, implying that $o(x^{i}y)=4$ for $1\leq i\leq 2m$. We can express every element in $Q_{4m}$ as $x^{i}y^{j}$ for $10\leq i\leq 2m-$ 1 and $0\leq j\leq 3$. Now, the element $x^{i}y^{j}$ is in the centre $Z(Q_{4m})$ if and only if it commutes with both $y$ and $x$. Clearly, $yx^{i}y^{j}=x^{i}y^{j}y$ if and only if $yx^{i}=x^{i}y$ if and only if $yx^{i}=yx^{-i}$, which is equivalent to $x^{2i}=1$. This forces $i\in\{0,m\}$. Thus, $x^{m}$ is in the centre $Z(Q_{4m})$. If $j\in\{0,2\}$ then $x^{i}y^{j}\in\{1,x^{m}\}$ which is contained in the centre $Z(Q_{4m})$. Consider the case $j\in\{1,3\}$. Then $x^{m}y^{j}x=x^{m+1}y^{j}$ if and only if $y^{j}x=xy^{j}$, which is equivalent $y^{j}x=y^{j}x^{-1}$ as $xy^{j}=y^{j}x^{-1}$ since $j$ is odd. But that implies $x^{2}=1$ which is a contradiction as $o(x)=2m>2$. Hence $Z(Q_{4m})=\{1,x^{m}\}$. $\qed$
The group of all permutations of $\{1,2,\ldots,n\}$ is the symmetric group $S_{n}$. For $n\geq 2$, its derived group is the alternating group $A_{n}$ consisting of all even permutations. The alternating group is simple for $n\geq 5$.
@@
be the Frobenius kernel of $G$. It is well-known (see, for example, [3, 35.24 and 35.25]) that $K\trianglelefteq G$, $G=K\rtimes H$, $C_{G}(h)\leq H$ for each $h\in H$, and $C_{G}(k)\leq K$ for each $k\in K$. Moreover, by the Thompson theorem on finite groups with fixed-point-free automorphisms of prime order [76, Theorem 1], $K$ is nilpotent. Frobenius groups are also characterised as transitive non-regular permutation groups in which non-identity elements fix at most one point.
The Frattini subgroup $\Phi(G)$ of a finite group $G$ is the intersection of all maximal proper subgroups of $G$. A $p$-group $G$ is extraspecial group $G$ if its centre $Z(G)$ is of order $p$ and coincides with its derived subgroup $G^{\prime}$ and the Frattini subgroup $\Phi(G)$ [45, Chapter 4, pp. 123]. Furthermore, there are exactly two non-isomorphic extraspecial groups of order $p^{2r+1}$, for each $r\geq 1$ and each prime $p$.
Every graph considered is simple: that is, it is undirected, with no loops and no multiple edges. Given a graph $\Gamma$, the vertex set and edge set of $\Gamma$ are denoted by $V(\Gamma)$ and $E(\Gamma)$, respectively. For a vertex subset $U\subset V(\Gamma)$ the induced subgraph $\Gamma[U]$ has vertex set $U$ and edge set $\{\{u,v\}\in E(\Gamma)\mid u,v\in U\}$. In other words, the subgraph induced by $U$ has vertex set $U$ and its edge set consists of all edges of $\Gamma$ between vertices in $U$.
Let $\mathcal{F}$ be a collection of finite graphs. If no induced subgraph of $\Gamma$ is isomorphic to a graph in $\mathcal{F}$, then $\Gamma$ is called $\mathcal{F}$-free; and if =$\mathcal{F}=\{\Delta\}$, we say that $\Gamma$ is $\Delta$-free. Sometimes we say instead that $\Gamma$ forbids $\mathcal{F}$ or $\Delta$. Denote by $P_{n}$ and $C_{n}$ the path of length $n-1$ and the cycle of length $n$, respectively. For distinct $x,y\in V(\Gamma)$, if $x$ is adjacent to $y$ in $\Gamma$, then we write $x\sim_{\Gamma}y$ to denote this edge $\{x,y\}$, or just by $x\sim y$ when $\Gamma$ is clear from the context. In particular, if $P_{n}$ is the subgraph of $\Gamma$ induced by the vertex subset $\{x_{1},x_{2},\ldots,x_{n}\}$ with adjacencies $x_{i}\sim x_{i+1},1\leq i\leq n-1$ we denote the induced path $P_{n}$ by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}.$
Similarly, in $\Gamma$, if the $n$-cycle $C_{n}$ is the subgraph induced by the subset of vertices $\{x_{1},x_{2},\ldots,x_{n}\}$, where $\{x_{1},x_{n}\}\in E(\Gamma)$ and $\{x_{i},x_{i+1}\}\in E(\Gamma)$ for all $11\leq i\leq n-$ 1, we denote it by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}\sim x_{1}.$